# EntroPE: Entropy-Guided Dynamic Patch Encoder for Time Series Forecasting

### v1

@@ 3.5 Global Transformer & Fusion Decoder
4 Experiments
Datasets. We evaluate EntroPE on seven widely-used long-term multivariate forecasting benchmarks: ETT family (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Electricity, and Exchange Rate datasets, covering diverse domains and temporal characteristics (detailed descriptions in Appendix Tab. 4).
Baselines. We benchmark against 14 state-of-the-art deep forecasting models from 2021–2025, organized by architecture. (1) Transformer-based models: iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021); (2) CNN-based model: TimesNet (Wu et al., 2023); (3) MLP-based models: TimeMixer (Wang et al., 2024), HDMixer (Huang et al., 2024), DLinear (Zeng et al., 2023); (4) Kernel-attention model: TimeKAN (Huang et al., 2025b); (5) Foundation and lightweight models: Time-FFM (Liu et al., 2024a), TimeBase (Huang et al., 2025a); (6) LLM-aligned and cross-modal models: LangTime (Niu et al., 2025), CALF (Liu et al., 2025); (7) Filter-based model: FilterTS (Wang et al., 2025). (Further details on baselines are discussed in Appendix A.1.)
Experimental Settings. Following standard practices in time series forecasting literature, we evaluate performance using Mean Squared Error (MSE) and Mean Absolute Error (MAE) metrics across multiple prediction horizons $T=\{96,192,336,720\}$ and fixed look-back window $L=96$. Adopting the same experiemnt setup as Huang et al. (2025b), baseline results up to 2024 are taken from TimeKAN, while other methods are re-evaluated using their official implementations. EntroPE experiemnts are conductd in 5 random seeds and we report the best results. (see Appendix A.3.)
4.1 Main Results
Table 1 summarizes average performance across all prediction horizons. EntroPE matches or surpasses, in many cases, recent variable-mixing, Transformer-based, CNN-based, MLP-based, foundation, lightweight, LLM-aligned, filter-based, and KAN-based models under the standard 96-step look-back with horizons up to 720. It retains PatchTST’s overall architecture but introduces two key innovations namely, (i) entropy-driven dynamic patching and (ii) adaptive embeddings. Relative to PatchTST, EntroPE achieves notable accuracy gains of approximately 20% on ETTh1, 15% on Electricity, and about 10% on average, while also reducing token count and improving efficiency by leveraging channel independence and non-overlapping patches. Beyond PatchTST, EntroPE consistently outperforms HDMixer’s length-extendable patching and improves upon TimeBase, though the latter remains highly efficient. On the Electricity dataset, iTransformer continues to offer the strongest performance-efficiency tradeoff given its channel-wise self-attention across 321 and 862 variables, respectively. CALF and LangTime also remain competitive, though their large model sizes (CALF has 18 million, LangTime has 500 million while EntroPE has 0.1 million range parameters) impose substantial costs. Across other benchmarks, EntroPE delivers robust improvements, underscoring the value of entropy-guided temporal segmentation for multivariate forecasting.
Table 1: Multivariate time series forecasting results on benchmark datasets. Results are averaged across prediction horizons $T=\{96,192,336,720\}$ with fixed input length $L=96$ for all datasets. Best results are highlighted in red and second-best in blue. A dash (-) indicates that the configuration was not found in the original implementation. (Full table: Appendix A.4)
Models ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Autoformer [2021] 0.496 0.487 0.450 0.459 0.588 0.517 0.327 0.371 0.338 0.382 0.227 0.338 0.613 0.539
FEDformer [2022] 0.498 0.484 0.437 0.449 0.448 0.452 0.305 0.349 0.309 0.360 0.214 0.327 0.519 0.429
DLinear [2023] 0.461 0.457 0.563 0.519 0.404 0.408 0.354 0.402 0.265 0.315 0.225 0.319 0.354 0.414
TimesNet [2023] 0.495 0.450 0.414 0.427 0.400 0.406 0.291 0.333 0.251 0.294 0.193 0.304 0.416 0.443
PatchTST [2023] 0.516 0.484 0.391 0.411 0.406 0.407 0.290 0.334 0.265 0.285 0.216 0.318 0.367 0.404
Time-FFM [2024] 0.442 0.434 0.382 0.406 0.399 0.402 0.286 0.332 0.270 0.288 0.216 0.299 0.338 0.391
HDMixer [2024] 0.448 0.437 0.384 0.407 0.396 0.402 0.286 0.331 0.253 0.285 0.205 0.295 - -
iTransformer [2024] 0.454 0.447 0.383 0.407 0.407 0.410 0.288 0.332 0.258 0.278 0.178 0.270 0.360 0.403
TimeMixer [2024] 0.459 0.444 0.390 0.409 0.382 0.397 0.279 0.324 0.245 0.276 0.182 0.272 0.378 0.410
TimeBase [2025] 0.463 0.429 0.409 0.425 0.431 0.420 0.290 0.332 0.252 0.279 0.227 0.296 - -
LangTime [2025] 0.437 0.425 0.375 0.392 0.397 0.392 0.284 0.321 0.252 0.273 0.201 0.285 0.361 0.400
TimeKAN [2025] 0.418 0.427 0.391 0.410 0.380 0.398 0.285 0.331 0.244 0.273 0.197 0.286 - -
FilterTS [2025] 0.440 0.432 0.375 0.399 0.386 0.397 0.279 0.323 0.253 0.280 0.184 0.275 0.355 0.400
CALF [2025] 0.441 0.435 0.372 0.395 0.396 0.391 0.280 0.321 0.250 0.274 0.177 0.266 - -
EntroPE [2025] 0.416 0.425 0.366 0.387 0.378 0.391 0.286 0.335 0.242 0.273 0.182 0.271 0.331 0.386
4.2 Ablation Studies
Model component analysis. To systematically evaluate each component’s contribution, we conduct comprehensive ablation studies across four datasets (ETTh1, ETTh2, ETTm1, Weather) using prediction horizons of 336 and 720 time steps, measuring performance via MSE. We progressively remove components from our full EntroPE architecture to create four configurations: (1) EntroPE (Full): Dynamic patching + Adaptive encoder + Fusion decoder; (2) EntroPE - Dynamic: Static patching + Adaptive encoder + Fusion decoder; (3) EntroPE - (Dynamic Patching + Adaptive Encoder): Static patching + Max pooling + Fusion decoder; (4) EntroPE - (Dynamic Patching + Adaptive Encoder + Fusion Decoder): Static patching + Max pooling + Flattened output.
Table 2: Ablation study showing component importance and dynamic vs. static patching comparison. Best results are highlighted in red and second-best in blue.
Dataset T Full -EDP -EDP -APE -EDP -APE -FD Baselines
Dynamic Static(1) Static(8) Static(16) Pool + FD Pool + Flat PatchTST iTrans.
ETTh1 336 0.429 0.425 0.441 0.444 0.438 0.519 0.501 0.487
720 0.439 0.460 0.469 0.477 0.461 0.527 0.503 0.500
ETTh2 336 0.355 0.428 0.435 0.439 0.439 0.462 0.427 0.428
720 0.397 0.448 0.470 0.463 0.466 0.456 0.436 0.427
ETTm1 336 0.393 0.401 0.409 0.415 0.402 0.427 0.421 0.426
720 0.445 0.451 0.459 0.452 0.457 0.483 0.462 0.491
Weather 336 0.258 0.261 0.265 0.270 0.262 0.262 0.284 0.278
720 0.341 0.343 0.353 0.359 0.347 0.346 0.356 0.358
Dynamic vs. Static Patching. In addition to the component ablations, we compare EntroPE with dynamic patching against three static (fixed-length, non-overlapping) patching schemes. Experiments are conducted on ETTh1, ETTm1, and Weather datasets with a fixed input length of 96 and forecasting horizons of 336 and 720. Dynamic patching achieves the lowest MSE in most settings (Table 2, while single-token input (patch length of 1) remains competitive, notably without leveraging the Dynamic Patching scheme. However, the computational cost of this setting is substantially higher, as the self-attention mechanism must attend to every time step individually. Figure 3 illustrates the efficiency advantage of dynamic patching over static alternatives under different configurations.
Figure 3: Ablation study for Dynamic and Static(patch_length) patching with single epoch duration.
[…]
Figure 4: Ablation study for threshold sensitivity with epoch duration.
@@ 4.3 Efficiency Analysis
Figure 5: EntroPE efficiency analysis.
As shown in Figures 3 and 4, our model adapts to different patching schemes, directly affecting computational efficiency. In the dynamic setting, the model maintains superior predictive performance while operating within an optimal computation budget. We evaluate efficiency on the ETTm1 dataset ($\rightarrow96$) by comparing recent models in terms of Multiply-Accumulate Operations (MACs), Mean Squared Error (MSE), and the number of trainable parameters. Figure 5 demonstrates that our model achieves a more favorable performance–efficiency trade-off, with bubble size proportional to the parameter count. Notably, while TimeMixer and TimeKAN occupy a nearby region in the trade-off space, our model attains a superior position, especially with its transformer-based architecture.
@@ 5 Conclusion
We introduced EntroPE, a temporally informed dynamic patch encoding framework that extends time series transformer architectures through entropy-guided boundary detection. Instead of partitioning sequences at arbitrary intervals, EntroPE strategically places patch boundaries at natural temporal transitions, preserving temporal coherence while supporting efficient variable-length patch processing. This design provides a principled way to incorporate information-theoretic signals into patching, yielding improved forecasting accuracy alongside practical computational benefits. Beyond its empirical performance, EntroPE highlights the importance of respecting the intrinsic structure of time series, opening new research directions in adaptive sequence modeling and context-aware temporal representation learning.
Appendix A Appendix
A.1 Baseline Details

### v2

@@ 4.3 Ablation Study
Figure 4: Threshold sensitivity with epoch duration.
@@ 4.4 Efficiency Analysis
As shown in Figures 3 and 4, our model adapts to different patching schemes, directly affecting computational efficiency. We evaluate efficiency on the ETTm1 dataset ($\rightarrow96$) by comparing recent models in terms of Multiply-Accumulate Operations (MACs), MSE, and the number of trainable parameters. Figure 5 demonstrates that our model achieves a more favorable performance-efficiency trade-off, with bubble size proportional to the parameter count. Notably, while TimeMixer and TimeKAN (Huang et al., 2025b) occupy a nearby region in the trade-off space, our model attains a superior position, especially with its transformer-based architecture. (See full table: Tab. 13)
Figure 5: EntroPE’s efficiency analysis on ETTm1 dataset.
@@ 5 Conclusion
We introduced EntroPE, a dynamic patch encoding framework that uses entropy-guided boundary detection to identify natural temporal transitions in time series. By placing patch boundaries at information-rich transitions rather than fixed intervals, EntroPE preserves temporal coherence while enabling efficient variable-length processing. This information-theoretic approach improves performance across downstream tasks with practical computational benefits. Beyond empirical gains, EntroPE demonstrates the value of respecting intrinsic temporal structure, opening new directions in adaptive context-aware sequence modeling.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Preliminaries & Motivation
A.1 Method Overview
Our approach consists of two distinct phases: (1) entropy model pre-training for uncertainty quantification, and (2) task-specific training with entropy-guided dynamic patching. Crucially, the entropy model parameters remain frozen during task-specific training, serving solely to provide patch boundaries.
A.2 Theoretical Motivation
We provide a theoretical justification for entropy-guided patch segmentation. Our framework rests on two principles: (i) conditional entropy characterizes irreducible predictive uncertainty, and (ii) segmenting sequences at high-entropy points yields patches with lower intra-segment uncertainty and improved temporal coherence.
Proposition A.1 (Conditional Entropy Lower-Bounds Irreducible Prediction Loss).
Let $\tau_{1:L}$ denote quantized time-series tokens and let $p_{\star}(\tau_{t+1}\mid\tau_{\leq t})$ be the true next-token conditional distribution. For any predictive model $q(\tau_{t+1}\mid\tau_{\leq t})$, the expected one-step negative log-likelihood satisfies
$\mathbb{E}_{p_{\star}}\!\left[-\log q(\tau_{t+1}\mid\tau_{\leq t})\right]\;\geq\;H_{\star}(\tau_{t+1}\mid\tau_{\leq t}),$ (17)
with equality if and only if $q(\cdot\mid\tau_{\leq t})=p_{\star}(\cdot\mid\tau_{\leq t})$ almost surely.
Proof sketch.
By the cross-entropy decomposition,
$\mathbb{E}_{p_{\star}}[-\log q]=H_{\star}(\tau_{t+1}\mid\tau_{\leq t})+\mathrm{KL}\!\left(p_{\star}(\cdot\mid\tau_{\leq t})\;\|\;q(\cdot\mid\tau_{\leq t})\right),$
and since $\mathrm{KL}(\cdot\|\cdot)\geq 0$, the bound follows, with equality if and only if $q=p_{\star}$ almost surely. ∎
Proposition A.2 (High-Entropy Boundaries Reduce Intra-Patch Predictive Uncertainty).
Let a segmentation $\mathcal{S}=\{1=b_{1}<b_{2}<\cdots<b_{P+1}=L+1\}$ partition a sequence into patches $b_{j},b_{j+1})$. Define the total within-patch uncertainty mass as
$U(\mathcal{S})=\sum_{j=1}^{P}\sum_{t=b_{j}}^{b_{j+1}-1}H_{\star}(\tau_{t+1}\mid\tau_{\leq t}).$ (18)
Consider a candidate boundary $b$ within a patch $[s,e)$, and suppose $H_{\star}(\tau_{b+1}\mid\tau_{\leq b})$ is a local maximum or exhibits a sharp positive increment relative to its neighborhood. Then introducing a boundary at $b$ yields a segmentation $\mathcal{S}^{\prime}$ such that the average within-patch entropy
$\frac{1}{|[b_{j},b_{j+1})|}\sum_{t=b_{j}}^{b_{j+1}-1}H_{\star}(\tau_{t+1}\mid\tau_{\leq t})$
is no greater in at least one resulting patch and strictly lower whenever $H_{\star}$ exhibits a concentrated spike at $b$.
Proof sketch.
Let $[s,e)$ contain $b\in(s,e)$. Splitting into $[s,b)$ and $[b,e)$ preserves total uncertainty mass, but redistributes it across two segments. If $H_{\star}(\tau_{b+1}\mid\tau_{\leq b})$ is locally elevated, then keeping $b$ interior inflates the mean uncertainty of the full segment. By relocating $b$ to a boundary, the high-entropy region is isolated to the segment edge, lowering the average entropy within at least one subsegment. This reduces expected intra-patch predictive uncertainty while preserving segmentation budget. ∎
Connection to EntroPE.
Proposition A.1 establishes conditional entropy as a fundamental lower bound on achievable predictive loss, implying that peaks in entropy correspond to irreducible uncertainty in temporal evolution. Proposition A.2 further motivates entropy-guided segmentation by showing that placing boundaries at high-entropy or sharply rising entropy points yields patches whose interiors exhibit lower average predictive uncertainty and greater temporal coherence.
EntroPE operationalizes this principle by training a lightweight causal transformer on quantized time series, freezing it, and computing per-timestep conditional entropy from the predicted token distribution. Since the true entropy $H_{\star}$ is unknown, we approximate it using the frozen model’s predictive distribution, $H_{t}\approx H_{\star}(\tau_{t+1}\mid\tau_{\leq t})$, and apply quantile-based thresholding (Eq. 56) to construct a boundary mask for dynamic patching. Our segmentation depends only on relative uncertainty structure, not absolute probability calibration.
Remark A.3 (Quantization Does Not Limit Predictions).
While we use quantization for entropy computation, the downstream prediction heads operate on continuous representations (Sec. 3.6). This decoupling eliminates range constraints: predictions are not limited to the discrete vocabulary $\mathcal{V}$ used for boundary detection.
This theoretical framework justifies our design: EntroPE identifies natural transition points where predictive uncertainty is high, ensuring patches are internally coherent while boundaries mark regime changes.
A.2.1 Phase I: Entropy Model Pre-training
Given a time series dataset $\mathcal{D}_{\text{pretrain}}=\{X^{(i)}\}_{i=1}^{N}$ where $X^{(i)}\in\mathbb{R}^{C\times L}$, we first quantize each continuous sequence into discrete tokens. For a univariate sequence $x=[x_{1},\ldots,x_{L}]$, we apply quantization $q:\mathbb{R}\to\mathcal{V}$ (defined in Eq. 38) to obtain tokenized sequence $\tau=[\tau_{1},\ldots,\tau_{L}]$ where $\tau_{t}=q(x_{t})\in\mathcal{V}$ and $|\mathcal{V}|=V$.
Automatic Quantization Range. To avoid manual threshold selection, we determine bin boundaries $\{b_{1},\ldots,b_{V-1}\}$ from training data statistics. Let $X_{\text{train}}$ denote z-score normalized training values. We compute empirical quantiles:
$q_{\text{low}}=Q_{\epsilon/2}(X_{\text{train}}),\quad q_{\text{high}}=Q_{1-\epsilon/2}(X_{\text{train}})$ (19)
where $\epsilon=0.005$ captures 99.5% of the distribution. The symmetric quantization radius is $R=\max(|q_{\text{low}}|,|q_{\text{high}}|)$, and the interval [R,R][-R,R] is uniformly divided into $V$ bins.
Model Training. We train a lightweight causal transformer $f_{\theta_{\text{ent}}}:\mathcal{V}^{t}\to\Delta^{V}$ (where $\Delta^{V}$ denotes the probability simplex over vocabulary $\mathcal{V}$) to perform next-token prediction. The model is optimized via:
$\theta_{\text{ent}}^{*}=\arg\min_{\theta_{\text{ent}}}\mathbb{E}_{X\sim\mathcal{D}_{\text{pretrain}}}\left[\mathcal{L}_{\text{CE}}(X;\theta_{\text{ent}})\right]$ (20)
where the cross-entropy loss for a sequence is:
$\mathcal{L}_{\text{CE}}(X;\theta_{\text{ent}})=-\frac{1}{L-1}\sum_{t=1}^{L-1}\log p_{\theta_{\text{ent}}}(\tau_{t+1}|\tau_{\leq t})$ (21)
Once converged, we freeze $\theta_{\text{ent}}^{*}$ and use it exclusively for entropy computation, not for prediction or token generation. This decoupling is critical: the entropy model learns predictive uncertainty patterns, while downstream task models operate on continuous representations.
A.2.2 Phase II: Entropy-Guided Dynamic Segmentation
Entropy Computation.
For a new input sequence $X\in\mathbb{R}^{C\times L}$, we process each channel independently. For channel $c$, we compute the quantized sequence $\tau^{(c)}=[q(x_{1}^{(c)}),\ldots,q(x_{L}^{(c)})]$ and use the frozen model $\theta_{\text{ent}}^{*}$ to calculate conditional entropy at each position:
$H(x_{t}^{(c)};\theta_{\text{ent}}^{*})=-\sum_{v\in\mathcal{V}}p_{\theta_{\text{ent}}^{*}}(\tau_{t+1}^{(c)}=v|\tau_{\leq t}^{(c)})\log p_{\theta_{\text{ent}}^{*}}(\tau_{t+1}^{(c)}=v|\tau_{\leq t}^{(c)})$ (22)
yielding entropy sequence $\mathbf{H}^{(c)}=[H(x_{1}^{(c)}),\ldots,H(x_{L}^{(c)})]\in\mathbb{R}^{L}$.
Theoretical Motivation. High conditional entropy $H(x_{t}|x_{<t})$ indicates temporal transitions where $x_{t}$ is less predictable from history. By placing boundaries at high-uncertainty regions, patches become internally coherent (low intra-patch entropy) while boundaries mark regime changes. Unlike fixed-length patching, our proposed method doesn’t ignore predictive structure.
Adaptive Threshold Determination.
Rather than fixed thresholds requiring dataset-specific tuning in general, we introduce sample-adaptive thresholds via quantile estimation. For the $\alpha$-th quantile (e.g., $\alpha=0.75$), we derive:
$\theta_{\text{abs}}^{(c)}$ $=Q_{\alpha}(\mathbf{H}^{(c)})\quad\text{(absolute entropy threshold)}$ (23)
$\Delta\mathbf{H}^{(c)}$ $=[H(x_{t}^{(c)})-H(x_{t-1}^{(c)})]_{t=2}^{L}\quad\text{(entropy differences)}$ (24)
$\gamma_{\text{rel}}^{(c)}$ $=Q_{\alpha}(\Delta\mathbf{H}^{(c)})\quad\text{(relative entropy threshold)}$ (25)
Boundary Detection.
We generate a binary boundary mask $M^{(c)}\in\{0,1\}^{L}$ where $M_{t}^{(c)}=1$ indicates a patch boundary at position $t$. Position $t$ becomes a boundary when:
$M_{t}^{(c)}=\mathbb{1}\left[H(x_{t}^{(c)})>\theta_{\text{abs}}^{(c)}\land\Delta H(x_{t}^{(c)})>\gamma_{\text{rel}}^{(c)}\land M_{t-1}^{(c)}=0\right]$ (26)
with $M_{1}^{(c)}=1$ (first position always starts a patch).
Patch Segmentation.
Given boundary mask $M^{(c)}$, we extract boundary positions $\mathcal{B}^{(c)}=\{t:M_{t}^{(c)}=1\}=\{b_{1}^{(c)},b_{2}^{(c)},\ldots,b_{P^{(c)}+1}^{(c)}\}$ where $b_{1}^{(c)}=1$ and $b_{P^{(c)}+1}^{(c)}=L+1$. The sequence is then partitioned into $P^{(c)}$ non-overlapping patches:
$x^{(c)}=\bigcup_{j=1}^{P^{(c)}}p_{j}^{(c)},\quad p_{j}^{(c)}=[x_{b_{j}^{(c)}},x_{b_{j}^{(c)}+1},\ldots,x_{b_{j+1}^{(c)}-1}]\in\mathbb{R}^{\ell_{j}^{(c)}}$ (27)
where $\ell_{j}^{(c)}=b_{j+1}^{(c)}-b_{j}^{(c)}$ is the length of patch $j$, and $\sum_{j=1}^{P^{(c)}}\ell_{j}^{(c)}=L$.
A.2.3 Phase III: Task-Specific Training with Fixed Boundaries
Patch Encoding.
Each variable-length patch $p_{j}^{(c)}$ is encoded into a fixed-dimensional representation via the Adaptive Patch Encoder (APE). First, we obtain time-point embeddings through an embedding layer $E:\mathcal{V}\to\mathbb{R}^{d_{t}}$:
$\mathbf{h}_{t}^{(c)}=E(q(x_{t}^{(c)})),\quad t=1,\ldots,L$ (28)
For patch $p_{j}^{(c)}$ spanning positions $\mathcal{T}_{j}^{(c)}=\{b_{j}^{(c)},b_{j}^{(c)}+1,\ldots,b_{j+1}^{(c)}-1\}$, we initialize its representation via max pooling:
$\mathbf{z}_{j}^{(c),0}=\text{MaxPool}\left(\{\mathbf{h}_{t}^{(c)}\}_{t\in\mathcal{T}_{j}^{(c)}}\right)\in\mathbb{R}^{d_{p}}$ (29)
We then refine patch embeddings through $N$ cross-attention layers where patch embeddings query only their constituent time points:
$\mathbf{z}_{j}^{(c),n}=\mathbf{z}_{j}^{(c),n-1}+W_{o}\sum_{i\in\mathcal{T}_{j}^{(c)}}\alpha_{ji}^{(n)}W_{v}(\mathbf{h}_{i}^{(c),n-1})$ (30)
where attention weights are computed only within the patch:
$\alpha_{ji}^{(n)}=\frac{\exp\left(\frac{[W_{q}(\mathbf{z}_{j}^{(c),n-1})]^{\top}W_{k}(\mathbf{h}_{i}^{(c),n-1})}{\sqrt{d_{k}}}\right)}{\sum_{i^{\prime}\in\mathcal{T}_{j}^{(c)}}\exp\left(\frac{[W_{q}(\mathbf{z}_{j}^{(c),n-1})]^{\top}W_{k}(\mathbf{h}_{i^{\prime}}^{(c),n-1})}{\sqrt{d_{k}}}\right)}$ (31)
This produces final patch embeddings $P^{(c)}=[\mathbf{z}_{1}^{(c),N},\ldots,\mathbf{z}_{P^{(c)}}^{(c),N}]\in\mathbb{R}^{P^{(c)}\times d_{p}}$ and encoder hidden states $H^{(c)}=[\mathbf{h}_{1}^{(c),N},\ldots,\mathbf{h}_{L}^{(c),N}]\in\mathbb{R}^{L\times d_{t}}$.
Batch Processing with Variable Patches.
For a batch of $B$ samples with varying patch counts $\{P^{(b,c)}\}$, let $P_{\max}=\max_{b,c}P^{(b,c)}$. We construct a padded patch embedding tensor $\tilde{P}^{in}\in\mathbb{R}^{(B\cdot C)\times P_{\max}\times d_{p}}$ with attention mask $\mathcal{M}\in\{0,1\}^{(B\cdot C)\times P_{\max}}$ where $\mathcal{M}_{bc,j}=\mathbb{1}[j\leq P^{(b,c)}]$ to mask padded positions during self-attention.
Global Transformer.
Patch embeddings are processed through a standard transformer to capture inter-patch dependencies:
$\tilde{P}^{(c)}=\text{Transformer}(P^{(c)},\mathcal{M})\in\mathbb{R}^{P^{(c)}\times d_{p}}$ (32)
Fusion Decoder.
We fuse global patch context with fine-grained temporal information via cross-attention where time-point embeddings query patch representations:
$\tilde{H}^{(c)}=H^{(c)}+W_{o}\left(\text{softmax}\left(\frac{[W_{q}(H^{(c)})][W_{k}(\tilde{P}^{(c)})]^{\top}}{\sqrt{d_{k}}}\right)W_{v}(\tilde{P}^{(c)})\right)\in\mathbb{R}^{L\times d_{t}}$ (33)
This mechanism enables knowledge transfer from high-level patch representations back to detailed time-point embeddings, preserving both local temporal patterns and global contextual dependencies.
Appendix B Task-Specific Linear Projection
Task-Specific Heads.
The enriched representations are projected to task-specific outputs:
Forecasting: Channel-wise linear projection
$\hat{Y}^{(c)}=g_{\text{fc}}\left(\text{Flatten}(\tilde{H}^{(c)})\right)\in\mathbb{R}^{T},\quad c=1,\ldots,C$ (34)
with MSE loss: $\mathcal{L}_{\text{forecast}}=\frac{1}{C\cdot T}\sum_{c=1}^{C}\sum_{t=1}^{T}(Y_{c,t}^{(c)}-\hat{Y}_{c,t}^{(c)})^{2}$
Classification: Flatten all channels and project to class logits
$\hat{y}=g_{\text{cls}}\left(\text{Flatten}([\tilde{H}^{(1)},\ldots,\tilde{H}^{(C)}])\right)\in\mathbb{R}^{K}$ (35)
with cross-entropy loss: $\mathcal{L}_{\text{class}}=-\log\frac{\exp(\hat{y}_{y})}{\sum_{k=1}^{K}\exp(\hat{y}_{k})}$ where $y\in\{1,\ldots,K\}$ is the true label.
Anomaly Detection: Channel-wise reconstruction
$\hat{X}^{(c)}=g_{\text{rec}}\left(\text{Flatten}(\tilde{H}^{(c)})\right)\in\mathbb{R}^{L},\quad c=1,\ldots,C$ (36)
with MSE loss: $\mathcal{L}_{\text{anomaly}}=\frac{1}{C\cdot L}\sum_{c=1}^{C}\sum_{t=1}^{L}(X_{c,t}^{(c)}-\hat{X}_{c,t}^{(c)})^{2}$
Training Protocol.
During task-specific training, the entropy model $\theta_{\text{ent}}^{*}$ remains frozen and is used only to compute patch boundaries via Eq. (8)-(10). Only the encoder $E$, APE, transformer, fusion decoder, and task head parameters $\Phi=\{E,\theta_{\text{APE}},\theta_{\text{trans}},\theta_{\text{FD}},\theta_{\text{task}}\}$ are optimized:
$\Phi^{*}=\arg\min_{\Phi}\mathbb{E}_{(X,Y)\sim\mathcal{D}_{\text{task}}}\left[\mathcal{L}_{\text{task}}(X,Y;\Phi\mid\theta_{\text{ent}}^{*})\right]$ (37)
where $\mathcal{D}_{\text{task}}$ is the task-specific dataset and $\mathcal{L}_{\text{task}}\in\{\mathcal{L}_{\text{forecast}},\mathcal{L}_{\text{class}},\mathcal{L}_{\text{anomaly}}\}$.
Computational Complexity.
Let $L$ be sequence length, $P$ the average number of patches, and $d$ the hidden dimension. EntroPE’s complexity is $\mathcal{O}(Ld+P^{2}d)$ where $P\ll L$, compared to $\mathcal{O}(L^{2}d)$ for point-wise models. The entropy threshold $\alpha$ provides explicit control: lower $\alpha$ → more boundaries → higher $P$ → more computation but finer granularity.
Appendix C Anomaly Detection
Datasets.
We further evaluate EntroPE on five standard anomaly detection benchmarks (SMD (Su et al., 2019), MSL(Hundman et al., 2018), SMAP (Hundman et al., 2018), SWaT (Mathur and Tippenhauer, 2016), PSM (Abdulaal et al., 2021)) by replacing the forecasting head with an anomaly scoring mechanism. Table 4 shows F1 scores compared to recent methods.
Appendix D Baselines
Baselines and Experimental Settings.
Baselines and Experimental Settings. We followed the same experimental settings and adopted the same baselines in TSLANet. These are TSLANet, GPT4TS, TimesNet, PatchTST, ETSformer (Woo et al., 2022), FEDformer, LightTS (Zhang et al., 2024), DLinear.
Table 4: Anomaly detection F1 scores (%) on benchmark datasets.
Dataset EntroPE TSLANet GPT4TS TimesNet PatchTST ETSformer FEDformer LightTS DLinear
SMD 87.1 87.91 86.89 84.61 84.62 83.13 85.08 82.53 77.10
MSL 85.3 83.32 82.45 81.84 78.70 85.03 78.57 78.95 84.88
SMAP 74.2 75.96 72.88 69.39 68.82 69.50 70.76 69.21 69.26
SWaT 94.3 92.80 94.23 93.02 85.72 84.91 93.19 93.33 87.52
PSM 97.7 97.73 97.13 97.34 96.08 91.76 97.23 97.15 93.55
Average 87.72 87.54 86.72 85.24 82.79 82.87 84.97 84.23 82.46
Results.
EntroPE attains an average F1 score of 87.72%, competitive with or better than TSLANet (87.54%), GPT4TS (86.72%), TimesNet (85.24%), and other baselines. These results confirm that the entropy-guided patching mechanism generalizes beyond forecasting to anomaly detection tasks.
D.1 Baseline Details
See Table 5 for the details about each baseline.